( how are they diffrent and what is it that its for)
Terrible question....
But heres the answer!468436845484

Answer by hehe(5)

distributive is kinda like this
>> Distributive of addition
2+2 = 2+2
>> Same with distribuative of multiplcation and so on...
>> assosiative property
3+2= 2+3
>>communtive property
(2+2) * 2 = 2 * (2+2)